M<PXQ=(8x-3)°,m<PXS=(10x+30)°,find the value of ‘x’ and m<QXS


1
Expert's answer
2021-11-18T17:44:58-0500

"10x+30=2(8x-3)"

"10x+30=16x-6"

"16x-10x=30+6"

"6x=36"

"x=6"


"<QXS=<PXQ"

"<QXS=(8\u00d76-3)"

"<QXS=48-3"

"<QXS=45\u00b0"